### v2

# Threshold-less and Flexibly Tunable Frequency Comb via Floquet Engineering

Abstract
Frequency combs have revolutionized communication, metrology and spectroscopy. Numerous efforts have been dedicated to developing integrated combs, predominantly relying on Pockels or Kerr mechanisms. In this work, we propose and demonstrate a new type of frequency comb—Floquet cavity frequency comb—that does not rely on intrinsic material non-linearity. By periodically modulating the resonance frequency of a cavity, a Floquet cavity with multiple equally spaced frequency components is created. The pump tone interacts with the pre-modulated cavity, generating the output frequency comb. This approach offers a flexible tuning range and operates in a threshold-less manner, obviating the need to overcome nonlinear initiation threshold. We implement this on a microwave cavity optomechanical system on-chip. Compared to Kerr optomechanical combs, this approach efficiently generates comb with pump signal far from the cavity’s intrinsic frequency, and the power required for detection is reduced by approximately a factor of $10^{6}$, providing a promising platform for frequency comb generation.
Introduction
Figure 1: Floquet cavity and frequency comb. a Examples of cavities with periodically modulated resonance frequencies. b Model of the Floquet cavity with the modulation frequency $\Omega$ and strength $A$. c Time evolution of the intracavity field amplitude in the Floquet cavity obtained via numerical simulation. For demonstration purposes, cavity dissipation is neglected and $A\approx 25\Omega$. d The relative strength in the frequency domain of the Floquet cavity modes $\omega_{0}+m\Omega$ (where $m$ is the index of the Floquet states). The red dotted line marks the intrinsic frequency $\omega_{0}$. e Schematics of the spectral density of Floquet cavity quasi-energy with increasing modulation strength $A$ (where $n$ is the Fock basis). f The schematic diagram of the generation of Floquet cavity frequency comb. By injecting the pump tone into the pre-modulated cavity, photons are emitted across multiple modes, resulting in a frequency comb.
Frequency combs, composed of discrete, equally spaced frequencies [1, 2], have contributed to advancements in optical communication [3, 4], precision metrology [5, 6], spectroscopy [7, 8], and atomic clock [9, 10, 11]. The compact on-chip frequency combs (microcombs) include microresonator-based combs [12, 13, 14] and electro-optic (EO) combs [15, 16, 17, 18, 19]. Microresonator combs relying on Kerr nonlinearity have been successfully demonstrated in silica [12], silicon nitride [20], silicon carbide [21], diamond [22], and lithium niobate (LN) [23]. Integrated EO combs utilize the Pockels effect, which have also recently been realized in thin-film LN devices [15, 16, 17] and lithium tantalate [19]. Despite these remarkable advances, EO and microresonator combs are facing limitations of their generation efficiency [15, 24, 25] or tuning precision [26, 20]. Therefore, emerging paradigms for frequency comb generation that do not rely on intrinsic medium nonlinearity are highly desired [27, 28, 29], as they offer promising avenues to overcome the aforementioned limitations in applications such as astrocombs [30, 31] and molecular fingerprinting [7, 32]. The frequency combs based on engineered nonlinearity could also offer a scalable and low-cost manner for emerging fields such as integrated photonic data links [33], quantum networks [34], and multiphoton entanglement [35, 36, 37]. Notably, Floquet engineering [38], which involves the periodic time-dependent modulation, has the capability to control and generate new multi-mode sidebands by tuning energy spectra [39, 40, 41], as well as to induce interactions between modes [42, 43]. These Floquet sidebands demonstrate that Floquet engineering is a promising although underexplored approach for frequency comb generation, without necessitating Kerr nonlinearity or Pockels effect.
In this work, we propose and demonstrate a new type of frequency comb, termed the Floquet cavity frequency comb. Specifically, by periodically modulating the resonance frequency of a cavity, we create a Floquet cavity with multiple equally spaced frequency components. Then, a single-frequency pump tone is added, interacting with this pre-modulated multi-mode structure to generate the output frequency comb. Additionally, the absolute positions of the comb teeth are referenced to the frequency of the pump tone, with the intensity approximately symmetrically distributed around the cavity’s intrinsic frequency. Furthermore, in contrast to Kerr or EO combs, which require precise frequency tuning to match the free spectral range, our approach is flexibly tunable, with the spacing and number of comb teeth determined by the modulation source. The frequency components of the Floquet cavity are predefined, enabling the frequency comb to exceptionally operate without the need to overcome a threshold condition.
Based on the proposed theoretical framework, such a frequency comb can be realized across a wide range of physical systems. Here, we demonstrate its implementation for modulating the resonance frequency of a superconducting microwave cavity via a mechanically compliant capacitor. To achieve sinusoidal modulation, we additionally employ an auxiliary cavity to establish a conventional optomechanical coupling, driving the mechanical oscillator from thermal state into self-induced oscillation [44]. In contrast to Kerr optomechanical combs [45, 46, 47], which rely on the nonlinear coupling between phonons and photons [48, 49, 50], our approach achieves a significantly extended tuning range for the pump signal, including frequencies that are substantially red-detuned from the cavity’s natural resonance. Furthermore, this approach reduces the input power required for detection by a factor of approximately $10^{6}$, which is limited primarily by the detection efficiency. This positions Floquet cavities as a promising platform for frequency comb generation.
Results
Floquet cavity frequency comb
We introduce a general Floquet cavity, defined as a cavity whose resonance frequency $\omega_{\mathrm{c}}$ is periodically modulated in time. It can be mapped to counterparts in well-established physical systems, including Fabry–Pérot resonators of movable mirror or microwave circuits with adjustable capacitance or inductance, as illustrated in Fig. 1a. The resonance modes of these Floquet cavities transform from a single-frequency state to a superposition of multiple frequencies (see Supplementary Note I.A). The Hamiltonian of the Floquet cavity is given by $H_{\mathrm{Floquet}}=\mathrm{\hbar}\left(\omega_{\mathrm{0}}+A\cos(\Omega t+\phi)\right)\hat{a}^{\dagger}\hat{a}$, where $\omega_{\mathrm{0}}$ is the intrinsic frequency, $A$ is the modulation strength, $\Omega$ is the modulation frequency, and we drop an irrelevant phase $\phi$, as illustrated in Fig. 1b.
In terms of Floquet theory, such periodic modulation of the cavity frequency effectively creates “Floquet sidebands” around the cavity’s intrinsic frequency. The quasi-energy of this Floquet cavity satisfies $\epsilon_{nm}/\hbar=n\omega_{\mathrm{0}}+m\Omega$, with the eigen-frequencies $\omega_{\mathrm{0}},\omega_{\mathrm{0}}\pm\Omega,\omega_{\mathrm{0}}\pm 2\Omega,\dots$(see Methods), where integer $n$ is the Fock basis of bare cavity and $m$ is the Floquet index. We present a numerical example illustrating the time evolution of the intracavity field in the Floquet cavity in Fig. 1c. A set of equally spaced components corresponding to the intracavity field in the Fourier domain is shown in Fig. 1d. The relative strength of the Floquet modes is determined by $J_{\mathrm{m}}(\beta)$, the Bessel function of the first kind with order $m$ (see Supplementary Note I.B and I.C). Here we define the dimensionless modulation coefficient $\beta=\frac{A}{\Omega}$. With a sufficiently large modulation strength, a substantial number of detectable Floquet modes can be produced.
Remarkably, since these modes are induced by Floquet engineering, they are not independent (as in modes in spatially separate cavities) but are correlated. The modulation strength $A/2$ determines the coupling strength between the different Floquet modes, thus forming what is referred to as a giant-mode cavity. A stronger modulation strength clearly enhances the generation of additional sidebands with appreciable intensity, as shown in Fig. 1e. Furthermore, this coupling enables energy input into the cavity to be transferred to different sidebands. Consequently, any input pump tone, even in the few-photon regime, is emitted across multiple sideband modes, forming a frequency comb with spacing $\Omega$, as illustrated in Fig. 1f.
Considering a input pump tone at the frequency $\omega_{\mathrm{in}}$, from the equation of motion, we obtain that the complex amplitude in the Floquet cavity satisfies (Supplementary Note II.A):
$a(t)=$ $\sqrt{\kappa_{\mathrm{e}}}a_{\mathrm{in}}e^{-i\Delta t}$ (1)
×$\times\sum_{m=-\infty}^{\infty}\sum_{l=-\infty}^{\infty}\frac{J_{l}(\beta)J_{m-l}(-\beta)}{il\Omega+\kappa/2-i\Delta}e^{im\Omega t},$
where $\Delta=\omega_{\text{in}}-\omega_{0}$ arises from the rotation of the frame, $\kappa$ represents the total dissipation of the cavity, $\kappa_{\mathrm{e}}$ denotes the external dissipation due to the coupling between the cavity and the external input/output, and $a_{\mathrm{in}}$ represents the externally injected pump field.
For a Floquet cavity, the modulation coefficient $\beta$ is predetermined and independent of $a_{\mathrm{in}}$, enabling threshold-less frequency comb operation, as indicated by Eq. (1). In other terms, the pump tone does not induce nonlinear effects but instead serves as a probe for a system with a prescribed time-varying response function. Furthermore, we found that the pump tone exhibits highly flexible tuning capabilities within the approximate range $-2\beta\Omega\lesssim\Delta\lesssim 2\beta\Omega$, as detailed in the Supplementary Note II.A. When $\Delta\neq 0$, the pump tone alters the relative intensities of the frequency comb components to some extent. In contrast, in Kerr optomechanical combs, the generation of the frequency comb by optomechanical non-linearity depends on the pump tone itself, meaning $\beta$ is influenced by both $a_{\mathrm{in}}$ and $\Delta$. Consequently, $a_{\mathrm{in}}$ must reach a sufficient intensity to overcome an initiation threshold. Additionally, $\Delta$ needs to be constrained within a range that can effectively excite the oscillator (For $\kappa<\Omega$, $\Delta\approx\Omega$) in Kerr optomechanical combs.
Figure 2: Construct the Floquet cavity in optomechanical systems. a Micrographs of the device, showing the entire chip (top) and zoomed-in drums (bottom). We select the larger drum (left) for the demonstration, while the other one is ignored. b Equivalent circuit model: straight arrows indicate the corresponding applied signals (blue for cavity-1, purple for cavity-2). c Cavity-1 serves as the auxiliary driving cavity to drive the mechanical oscillator into self-induced oscillation. Cavity-2 is modulated by the oscillator to construct the Floquet giant-mode cavity. d Mechanical oscillator spectrum transitioning from thermal state to self-induced oscillation regime; the red curve marks the spectrum at the critical drive power $P_{\mathrm{dc}}$. e The reflection spectrum of the cavity-2 measured by S21 as a function of increasing drive power. Colored dots: measurements; black lines: fits based on the reflection coefficient $R$. $X_{\mathrm{0}}$ is incorporated into $\omega_{\mathrm{c2}}$. The inset marks the driving frequency schematics.
The Device
We demonstrate Floquet cavity frequency comb in a superconducting microwave circuit, as illustrated in Fig. 2a. The resonance frequency of the cavity is modulated by a displacement-dependent capacitor. The sample used in this study is similar to that reported in our previous work [51, 52] (detailed fabrication in Supplementary Note III.A). Our device consists of two aluminum drumhead mechanical oscillators [53, 54] optomechanically coupled to two microwave cavities, as schematically shown in Fig. 2b. Here, we ignore one of the mechanical oscillators. The Hamiltonian of such a system can be expressed as
$H_{\mathrm{sys}}=\sum_{j=1,2}\hbar(\omega_{\mathrm{c},j}+\frac{g_{\mathrm{0},j}}{x_{\mathrm{zpf}}}\hat{x})\hat{a}_{j}^{\dagger}\hat{a}^{\mathstrut}_{j}+\hbar\omega_{\mathrm{m}}\hat{b}^{\dagger}\hat{b},$ (2)
where $\omega_{\mathrm{c},j}$ represents the $j$-th cavity’s intrinsic frequency. $\omega_{\mathrm{m}}$ represents the mechanical resonant frequency, and $a_{j}$ ($a_{j}^{\dagger}$) and $b$ ($b^{\dagger}$) represent the annihilation (creation) operators of the cavities and mechanical oscillator, respectively. $\hat{x}$ is the displacement of the mechanical oscillator, and $x_{\mathrm{zpf}}$ represents the zero-point fluctuation of the mechanical oscillator’s displacement. Besides, the coefficient $g_{\mathrm{0},j}$ represents the single photon coupling strength between the $j$-th cavity and mechanical oscillator.
Our entire device operates in a dilution refrigerator at a temperature of 10 mK (detailed setup in Supplementary Note III.B). The first cavity’s (cavity-1) intrinsic frequency is $\omega_{\mathrm{c}1}=2\pi\times 4.91$ GHz, with external decay rate $\kappa_{\mathrm{e}1}=2\pi\times 49.6$ kHz and intrinsic decay rate $\kappa_{\mathrm{i}1}=2\pi\times 298.9$ kHz. For the other microwave cavity (cavity-2), the intrinsic frequency is $\omega_{\mathrm{c}2}=2\pi\times 6.47$ GHz, accompanied by $\kappa_{\mathrm{e}2}=2\pi\times 21.7$ kHz and $\kappa_{\mathrm{i}2}=2\pi\times 245.7$ kHz. The oscillator is characterized to have frequency of $\omega_{\mathrm{m}}=2\pi\times 9.1$ MHz and the damping rate $\gamma_{\mathrm{m}}=2\pi\times 124$ Hz. Additionally, from the measurements based on frequency modulation technique, the single-photon coupling strengths are measured to be $g_{01}=2\pi\times 79$ Hz and $g_{02}=2\pi\times 46$ Hz. The detailed characterization method is described in Supplementary Note III.C.
Construct the Floquet cavity
We choose the cavity-2 and the mechanical oscillator together to form the Floquet cavity system with modulation frequency $\omega_{\mathrm{m}}$, as schematically shown in Fig. 2c. It is worth noting that, in the absence of excitation, the mechanical oscillator is in a thermal state in equilibrium, with an extremely small time-averaged amplitude. To achieve Floquet modulation of the cavity-2 in a sinusoidal periodic manner, we designate the cavity-1 as auxiliary driving cavity to drive the mechanical oscillator into a self-induced oscillation or phonon lasing state (Supplementary Note IV) through optomechanical parametric excitation [55, 56, 44].
We apply a blue-sideband drive at the frequency $\omega_{\mathrm{d}}=\omega_{\mathrm{c1}}+\omega_{\mathrm{m}}$ with drive power $P_{\mathrm{d}}$. The mechanical oscillator’s vibrations are gradually amplified and enter self-induced oscillation at a critical power $(P_{\mathrm{dc}})$ of $3\,\mathrm{dBm}$, as shown in Fig. 2d. Note that the power values correspond to the signal generator output powers and are reported in arbitrary units. We assume that the displacement is a classical state $x(t)=X_{\mathrm{m}}\cos(\omega_{\mathrm{m}}t+\phi_{\mathrm{m}})+X_{\mathrm{0}}$ [57]. Here, $X_{\mathrm{m}}$ is the classical amplitude and the $X_{\mathrm{0}}$ is the oscillation offset that is incorporated into $\omega_{\mathrm{c2}}$ (see Methods). Correspondingly, the modulation strength $A=\frac{g_{\mathrm{02}}X_{\mathrm{m}}}{x_{\mathrm{zpf}}}$ is controlled by the drive power $P_{\mathrm{d}}$.
We characterize the Floquet sidebands of cavity-2 by using a vector network analyzer (VNA) to extract the reflection coefficient $S_{\mathrm{21}}=R=1-\sqrt{\kappa_{\mathrm{e}}}\frac{\langle a(t)\rangle}{\langle a_{\mathrm{in}}(t)\rangle}$. Combining with Eq. (1), the depth of the different sidebands is also determined by the Bessel function (detailed in Supplementary Note II.B), as shown in Fig. 2e. These sidebands are symmetrically distributed with a spacing corresponding to the oscillator frequency $\omega_{\mathrm{m}}$. We studied the effect of drive power, the results are shown in Fig. 2e for $P_{\mathrm{d}}$ at -5 dBm, 3 dBm, and 10 dBm. Before the critical power point $P_{\mathrm{dc}}$, the mechanical oscillator stays in thermal state and thus only the intrinsic resonance frequency of cavity-2 can be detected. When $P_{\mathrm{d}}>P_{\mathrm{dc}}$, the self-induced mechanical oscillation modulates cavity-2 such that it can be effectively described as the Floquet cavity. Increasing the modulation strength enhances the ability of photons to transition to higher-order Floquet energy levels. Consequently, the intensity of the intrinsic frequency signal diminishes, while higher-order sidebands become more pronounced. The result from the free parameter fitting indicates that, under a 10 dBm drive, $\beta$ is approximately 1.92.
Emergence and evolution of Floquet comb
Figure 3: Emergence and evolution of Floquet comb. a Illustration of the setup. The blue arrow represents a fixed drive signal at $\Delta_{\mathrm{d}}=\omega_{\mathrm{m}}$ to excite mechanical oscillator. The red arrow in the Floquet cavity (cavity-2) represents the pump signal responsible for the generation of the frequency comb. b Emergence and evolution of frequency combs. Spectral responses are sorted into three types, and are marked in different colors as dots. We consider the frequency comb to be effective when the number of comb teeth $\geq 4$. The gray points represent the absence of the comb. The Floquet comb manifests in the regions marked with purple dots. In red-dotted region, cavity-2 itself produces a Kerr frequency comb. The black and red dashed lines are the dividing lines between the different regions. c Several examples of frequency comb corresponding to the dark points highlighted in b are shown, with each comb tooth marked by a line and a dot.
To achieve a stable self-induced oscillation, we applied a fixed drive power $P_{\mathrm{d}}=10$ dBm. As shown in Fig. 3a, we send a pump tone to cavity-2 and characterize the emergence and evolution of the frequency comb in the cavity response signal with different detuning $\Delta_{\mathrm{p}}=\omega_{\mathrm{p}}-\omega_{\mathrm{c2}}$ and power $P_{\mathrm{p}}$ (see Fig. 3b). The local minima occurring at the sideband frequencies $\Delta_{\mathrm{p}}=m\omega_{\mathrm{m}}$ are due to the fact that the pump frequency matches one of the Floquet sidebands, thereby allowing efficient photon injection. We demonstrate the flexible tuning capability of the Floquet frequency comb within the range of $-3\omega_{\mathrm{m}}\leq\Delta_{\mathrm{p}}\leq 3\omega_{\mathrm{m}}$. Notably, frequency comb can also emerge in the red sideband region of the cavity-2 ($\Delta_{\mathrm{p}}<0$), which is completely absent in Kerr optomechanical combs (for $\kappa<\Omega$) [47, 48, 49, 50, 57]. In theory, the pump power could be much lower, but it is constrained by the limitations of the detection accuracy in our experimental setup. Fig. 3c shows an example of the frequency comb at $\Delta_{\mathrm{p}}=-1\times\omega_{\mathrm{m}}$. As the pump power increases, the number of detectable frequency comb teeth increases, but the pump power itself does not affect the emergence of the frequency comb.
For comparison, we also independently characterized the emergence and evolution of the Kerr optomechanical frequency comb in cavity-2. As shown by the red data points in Fig. 3b, the minimum threshold power $P_{\mathrm{p-Kerr}}=11\ \mathrm{dBm}$ and occurs at the blue sideband $\Delta_{\mathrm{p}}\approx\Omega_{\mathrm{m}}$ because of maximized antidamping, triggering the optomechanical instability most efficiently.
Focusing on the Floquet mode of $m=1$ in Fig. 3b, the minimum generated power for the Floquet frequency comb is approximately reduced by six orders of magnitude compared to Kerr optomechanical combs. Moreover, for the $m=1$ mode, the width of the local minimum is significantly broader and shifts toward $m=0$. This occurs because, at this point, the pump not only serves as the probe of Floquet frequency comb, but also drives the mechanical oscillator acting similarly to the mechanism in Kerr optomechanical combs. The interaction between the pump signal and the mechanical oscillator results in a broader local minimum and the shift of this mode.
Effect of modulation strength
Figure 4: a Effect of drive and pump power on Floquet frequency comb generation at $m=-1,0,+1$(Red, purple, blue). The colored regions indicate the Floquet comb is generated, whereas the white areas represent the absence of a detectable frequency comb. The green dashed line represents the power range selected for subsequent analysis. b Example frequency comb for $m=-1$: frequency domain, time domain signals (Red: real part; Blue: imaginary part) and phase space trajectories. c Effect of modulation strength on the relative magnitudes of the frequency comb teeth, with error bars representing standard deviations. The gradient-colored points represent the peak values of the corresponding $m$-th comb tooth.
We investigate the impact of modulation strength on the Floquet frequency comb by setting the pump tone at $\Delta_{\mathrm{p}}/\omega_{\mathrm{m}}=-1,0,+1$ to match the $m$-th Floquet mode. As shown in Fig. 4a, while varying both the drive power and pump power, for $\Delta_{\mathrm{p}}=-\omega_{\mathrm{m}}$, the frequency comb is generated only when the Floquet modulation is activated, $P_{\mathrm{d}}>P_{\mathrm{dc}}$. The fluctuations in the minimum power value are attributed to measurement uncertainties. Notably, for $\Delta_{\mathrm{p}}=0$, when the pump power is sufficiently high, the frequency comb is generated before $P_{\mathrm{dc}}$. This is the result of the cooperative action of the optomechanical frequency comb under the external driving of two cavities. The results for $\Delta_{\mathrm{p}}=\omega_{\mathrm{m}}$ exhibit similar behavior. As at sufficiently high pump power, cavity-2 can independently generate a frequency comb via Kerr optomechanical interactions. Thus, we observe a gradual onset rather than a sharp transition with $\Delta_{\mathrm{p}}=\omega_{\mathrm{m}}$.
To minimize interference from the Kerr optomechanical frequency comb in our analysis, we focus on the pump at $\Delta_{\mathrm{p}}=-\omega_{m}$. Fig. 4b shows an example of Floquet frequency comb with $P_{\mathrm{d}}=11\,\mathrm{dBm}$ and $P_{\mathrm{p}}=0\,\mathrm{dBm}$. Time-domain signals and phase space trajectories indicate that the output comb maintains excellent coherence. The periodicity of the time-domain signal results from a 2 MHz offset between the demodulation frequency and the strongest peak, demonstrating the stability and coherence of the Floquet frequency comb in this regime. For the $\Delta_{\mathrm{p}}=-\omega_{m}$ detection, the resulting frequency comb intensity is not perfectly symmetric, which is attributed to the non-zero $\Delta$ term and $X_{\mathrm{0}}$ (see Supplementary Note V).
We further demonstrate the effect of modulation strength for the $\Delta_{\mathrm{p}}=-\omega_{m}$ case with $P_{\mathrm{p}}=0\,\mathrm{dBm}$, shown in Fig. 4c. As the modulation strength increases, the Floquet cavity’s ability to scatter photons into higher-order sidebands strengthens, resulting in a decrease in the intensity of the peak at $m=0$ and an increase in the intensities of peaks at $m=\pm 1,\pm 2,\dots$. With further increase in modulation strength, the intensity of the lower-order sidebands stabilizes and begins to decrease, while the effect of higher-order sidebands becomes more pronounced. At higher $P_{\mathrm{d}}$ values, it is noteworthy that the intensity of $m=-1$ exhibits an anomalous increase. This deviation is due to the $m=-1$ mode overlapping with the pump signal, preventing it from following the expected intensity decrease in sync with $m=+1$. Further details of the frequency combs and numerical simulations can be found in Supplementary Note V.
Discussion
We have proposed and demonstrated a new type of frequency comb via Floquet engineering. The approach developed in this work not only offers a threshold-less route for frequency comb emission but also significantly broadens the frequency tuning range. This advancement paves the way for the efficient generation and control of frequency combs, making them more suitable for low-power regimes required by sensitive systems. For example, in superconducting quantum computing devices, components such as superconducting resonators and qubits cannot operate under high photon inputs and cryogenic amplifiers have typically low saturation powers. Moreover, the generation of Floquet frequency comb is not limited to microwave optomechanical systems. Due to the extent to which the mechanical oscillator is driven, our demonstration shows up to approximately ten characteristic comb teeth. By optimizing stronger modulation sources, the number of frequency comb teeth can be rapidly expanded, with potential for self-referencing. Floquet frequency combs can be realized and applied in various systems exhibiting periodic modulation, such as SQUIDs [58], trapped ions [29], time crystals [59], and molecular or atomic systems [60], particularly in the optical domain. This holds promise for extending the capabilities of frequency combs in high dimensional quantum states [61], entanglement generation [36, 37], and integrated multi-frequency signal sources across more systems [58, 29].
Methods
Floquet cavity
Considering the Floquet Cavity Hamiltonian, the general form of the steady-state solution $\left|\psi_{\alpha},t\right\rangle$ to the Schrödinger equation can be written as
$\left|\psi_{\alpha},t\right\rangle=e^{-i\varepsilon_{\alpha}t}\left|u_{\alpha},t\right\rangle,\left|u_{\alpha},t\right\rangle=\left|u_{\alpha},t+T\right\rangle,$ (3)
Here, $\varepsilon_{\alpha}$ is the quasi-energy. So the Hamiltonian can be represented as a diagonal matrix $H(t)=\sum_{n}h_{n}(t)|n\rangle\langle n|$ under the Fock basis $\{|n\rangle\}$, with matrix elements of $h_{n}(t)=n\left(\omega_{\mathrm{0}}+A\cos\Omega t\right)$. Thus, the Schrödinger equation can be simplified as
$i\partial_{t}\psi_{n}(t)=n\left(\omega_{0}+A\cos\Omega t\right)\psi_{n}(t),$ (4)
here $\psi_{n}(t)=\langle n|\psi,t\rangle$. Combined with Eq. (3), the steady-state solution can be expressed as
$\left|\psi_{n},t\right\rangle$ $=e^{-in\omega_{0}t}\left|u_{n},t\right\rangle,$ (5)
$\left|u_{n},t\right\rangle$ $=e^{-in\Omega A\sin\Omega t}|n\rangle.$
We define $\left|u_{nm},t\right\rangle=\left|u_{n},t\right\rangle e^{im\Omega t}=u_{nm}(t)|n\rangle$, with $u_{nm}(t)=e^{-in\Omega A\sin\Omega t}e^{im\Omega t}$, which satisfies
$H_{F}\left|u_{nm},t\right\rangle=\varepsilon_{nm}\left|u_{nm},t\right\rangle.$ (6)
At this point, we have identified Brillouin regions over the frequency domain, where the single Brillouin region have a width of $\Omega$, and the different Brillouin regions are represented by the index $m$, called the Floquet index. And where $H_{F}=H(t)-i\partial_{t}$ is called the Floquet Hamiltonian, $\left|u_{nm},t\right\rangle$ is the time-periodic Floquet state, and quasi-energy $\varepsilon_{nm}=n\omega_{0}+m\Omega$. The $\left|u_{n},t\right\rangle$ can also be written as $\left|u_{n0},t\right\rangle$. Under the complete basis vector set $\left.\left\{\left|n_{mn}\right\rangle\right\rangle\right\}$ with $\left|n_{m}\right\rangle=|n\rangle e^{im\Omega t}$, the matrix elements of the Floquet Hamiltonian are
$\left.\left\langle\left\langle n_{m}\right|H_{F}\mid n_{m^{\prime}}^{\prime}\right\rangle\right\rangle=\langle n|H^{\left[m-m^{\prime}\right]}\left|n^{\prime}\right\rangle+m^{\prime}\Omega\delta_{mm^{\prime}}\delta_{nn^{\prime}},$ (7)
where $H^{[l]}$ is the Fourier component of $H(t)$. Since
$H(t)$ $=\sum_{n}n\left(\omega_{0}+A\cos\Omega t\right)|n\rangle\langle n|$ (8)
$=\sum_{l}H^{[l]}e^{il\Omega t},$
where $H^{[0]}=\sum_{n}n\omega_{c}|n\rangle\langle n|$, $H^{[\pm 1]}=\sum_{n}n\frac{A}{2}|n\rangle\langle n|$, and $H^{[l]}=0$ for $|l|>1$. Thus, the Eq. (7) can be further expressed as
$\left.\left\langle\left\langle n_{m}\right|H_{F}\mid n_{m^{\prime}}^{\prime}\right\rangle\right\rangle$ $=\delta_{nn^{\prime}}\left(n\omega_{0}\delta_{mm^{\prime}}+n\frac{A}{2}\delta_{m,m^{\prime}\pm 1}\right)$ (9)
$+m^{\prime}\Omega\delta_{mm^{\prime}}\delta_{nn^{\prime}},$
Obviously, $H_{F}$ is a block diagonal form about $n$, and we write the projection of $H_{F}$ in subspace $\left.\left\{\left|n_{m}\right\rangle\right\rangle\right\}$ with fixed $n$ as $H_{F}^{(n)}$ :
$H_{F}^{(n)}=\left[\begin{array}[]{ccccc}\ddots&\vdots&\vdots&\vdots&\begin{sideways}$\ddots$\end{sideways}\\ \cdots&n\omega_{0}-\Omega&n\frac{A}{2}&0&\cdots\\ \cdots&n\frac{A}{2}&n\omega_{0}&n\frac{A}{2}&\cdots\\ \cdots&0&n\frac{A}{2}&n\omega_{0}+\Omega&\cdots\\ \begin{sideways}$\ddots$\end{sideways}&\vdots&\vdots&\vdots&\ddots\end{array}\right]\text{, }$ (10)
which satisfies $H_{F}^{(n)}\left|u_{nm}\right\rangle\rangle=\varepsilon_{nm}\left|u_{nm}\right\rangle\rangle$ according to the Eq. (6). It can be observed that each Fock state energy level splits into $12m+$ 1 sublevels, consistent with the periodic nature of the Floquet Hamiltonian. The off-diagonal elements, proportional to $A/2$, characterize the coupling strength between adjacent Floquet states, thereby determining the transition rates between different energy levels. This coupling reflects the influence of the periodic driving field and results in transitions mediated by the exchange of discrete energy quanta of size $\Omega$. For more detailed derivations, please refer to the Supplementary Note I.B.
Self-induced oscillation by blue-detuned driving
In the blue-detuned regime, the optomechanical damping rate $\gamma_{\mathrm{opt}}$ is negative. Initially, this results in an increase in the oscillator’s effective temperature. As the overall damping rate $\gamma_{\mathrm{m}}+\gamma_{\mathrm{opt}}$ becomes negative, any small initial fluctuation will grow exponentially over time, eventually reaching a steady-state regime. These are referred to as self-induced optomechanical oscillations (or phonon lasing regime), where the mechanical oscillation amplitude stabilizes at a fixed value $X_{\mathrm{m}}$. The governing dynamical equations are given by
$\dot{a}_{1}(t)=\left[-\frac{\kappa_{1}}{2}+i\left(\Delta_{\mathrm{d}}-\frac{g_{\mathrm{01}}}{x_{\mathrm{zpf}}}x\right)\right]a_{\mathrm{1}}(t)+\sqrt{\kappa_{\mathrm{e1}}}a_{\mathrm{d}},$ (11)
$\frac{\mathrm{d}^{2}x(t)}{\mathrm{d}t^{2}}+\gamma_{\mathrm{m}}\frac{\mathrm{~{}d}x(t)}{\mathrm{d}t}+\omega_{\mathrm{m}}^{2}x(t)=\frac{F(t)}{m_{\mathrm{eff}}},$ (12)
where the radiation-pressure force $F(t)=\frac{\hbar g_{\mathrm{01}}}{x_{\mathrm{zpf}}}|a_{\mathrm{1}}(t)|^{2}$, $\kappa_{1}$ is the total decay rate of the cavity-1, $a_{\mathrm{d}}$ represents the drive photons, and $m_{\mathrm{eff}}$ is the effective mass of the mechanical oscillator.
Ignoring in the initial phase $\phi_{\mathrm{m}}$, we have $x(t)=X_{\mathrm{m}}\cos(\omega_{\mathrm{m}}t)+X_{\mathrm{0}}$. In steady state, we define an amplitude-dependent effective optomechanical damping rate as
$\gamma_{\mathrm{opt}}=-\frac{\left\langle F\dot{x}\right\rangle}{m_{\mathrm{eff}}\left\langle\dot{x}^{2}\right\rangle},$ (13)
which represents the ratio of the time-averaged power input due to this force to the corresponding mechanical energy lost. The time-averaged radiation-pressure force determines the oscillation offset $X_{\mathrm{0}}$, and by combining the balance condition of the overall damping rate, we obtain
[…]
With the ansatz, the solution of $a_{\mathrm{1}}(t)$ in blue-sideband drive can be written in a Fourier series $a_{\mathrm{1}}(t)=e^{-i\varphi(t)}\sum_{k}a_{k}e^{ik\omega_{\mathrm{m}}t}$, with coefficients:
$a_{k}=\sqrt{\kappa_{\mathrm{e1}}}a_{\text{d}}\frac{J_{k}(g_{\mathrm{01}}X_{\mathrm{m}}/x_{\mathrm{zpf}}\omega_{\mathrm{m}})}{ik\omega_{\mathrm{m}}+\frac{\kappa}{2}-i(\omega_{\mathrm{m}}-g_{\mathrm{01}}X_{\mathrm{0}}/x_{\mathrm{zpf}})},$ (15)
and the global phase $\varphi(t)=(g_{\mathrm{01}}X_{\mathrm{m}}/x_{\mathrm{zpf}}\omega_{\mathrm{m}})\sin(\omega_{\mathrm{m}}t)$. By substituting this Fourier series into Eq. (14), we obtain the implicit equation for $X_{\mathrm{0}}$ and $X_{\mathrm{m}}$:
[…]
The series can be efficiently summed numerically to obtain the explicit dependence of $X_{\mathrm{m}}$ and $X_{\mathrm{0}}$ on the system parameters. Mathematically, the onset of small-amplitude oscillations, starting from $A=0$, represents a Hopf bifurcation. In this regime, $A\propto\sqrt{|a_{\mathrm{d}}|^{2}-|a_{\mathrm{d}}|_{\mathrm{th}}^{2}}$, where $|a_{\mathrm{d}}|_{\mathrm{th}}^{2}$ is the threshold value [55, 56, 47, 57]. For more detailed derivations, please refer to the Supplementary Note IV.
Data availability
All other data relevant to this study are available from the corresponding authors. Source data are provided with this paper.
Acknowledgments
We would like to express our gratitude for the fruitful discussions with Kaiye Shi, and Qiongyi He. Y. Liu acknowledges the support of Beijing Natural Science Foundation (Z240007), National Natural Science Foundation of China (Nos. 92365210 and 12374325), CAST Young Elite Scientists Sponsorship Program (Grant No. 2023QNRC001), and Beijing Municipal Science and Technology Commission (Grant No. Z221100002722011). This work was also supported by the Innovation Program for Quantum Science and Technology (2021ZD0302200), the Chinese Academy of Sciences (No. GJJSTD20200001), the National Key R&D Program of China (Grants No. 2021YFB3202800).
Author contributions
S. W. conducted the measurements, Y. L. conceived the project and coordinated the research. All authors contributed to the analysis of the measurement results, wrote the manuscript, and provided their final approval for publication.
Competing interests
The authors declare no competing interests.